OpenStudy (anonymous):

Please Help!! Suppose that the functions s and t are defined for all real numbers x as follows. s(x)=x^3 t(x)=4x^2 Write the expressions for (s+t)(x) and (s-t)(x) and evaluate (s*t)(-1).

geerky42 (geerky42):

HINT: \((s+t)(x) ~~=~~ s(x) + t(x)\), \((s-t)(x) ~~=~~ s(x) - t(x)\) and \((s\cdot t)(x) ~~=~~ s(x)\cdot t(x)\)